OpenStudy (micahm):

f(x)=3x+5 what is the range

OpenStudy (anonymous):

f(x) = 3x + 5 f(- 1) = - 3 + 5 = 2 f(0) = 5 f(1) = 8 f(2) = 11 f(3) = 14 g {2, 5, 8, 11, 14}

OpenStudy (micahm):

thank you